The sum of 2 numbers is 87. If twice the smaller number is subtracted from the larger number, the result is 12. Find the two numbers. How do I solve it step by step?
Answer by chosenpoint(26)

Linear-systems/249574 (2009-12-14 04:33:42): The sum of 2 numbers is 87. If twice the smaller number is subtracted from the larger number, the result is 12. Find the two numbers. How do I solve it step by step?

We need to set up a system of equations using x and y, where y is the smaller number.

x+y=87 given

x-2y=12 given

We can then easily subtract the bottom equation from the top equation in order to find the value of y, which leads to

(x-x)+y-(-2y)=87-12 subtract bottom equation from top equation

3y=75 simplify

y=25 divide both sides by 3

x+y=87 given

x+25=87 substitute 25 in for y

x=62 subtract 25 from both sides

Final Answer: 62 and 25

Answer 181745 by chosenpoint(19) on 2009-12-14 04:44:03